OpenStudy (anonymous):

Simplify. x^2-16/x^2-6x+9/x^2-3x-4/x^2-2x-3

OpenStudy (anonymous):

@Hero

OpenStudy (nikato):

is there any paranthesis?

OpenStudy (anonymous):

One sec

hero (hero):

I interpret it as this: \[\frac{x^2 - 16}{x^2 - 6x + 9} \div \frac{x^2 - 3x - 4}{x^2 - 2x -3}\]

OpenStudy (anonymous):

You are right

OpenStudy (nikato):

i think it';; be easier if u factor out everything first

hero (hero):

Which in that case, you can simply flip the second fraction change it to multiplication: \[\frac{x^2 - 16}{x^2 - 6x + 9} \times \frac{x^2 - 2x -3}{x^2 - 3x - 4}\]

hero (hero):

Next, you can factor each expression and then cancel factors of one to simplify.

OpenStudy (anonymous):

I don't get how to factor out the trinomials

OpenStudy (nikato):

do u know the box and diamond method?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I know that you are suppose to multiply the first one and second one then find one that adds up to the middle one

OpenStudy (nikato):

yea, its like that u see how the trinomials are written in the form Ax^2 + Bx +C and inthe pic, the top # is A times C and the bottom is just B ur tryin to find 2 factors of AC that add up to B we can try factorring x^2 -2x -3

OpenStudy (nikato):

 so what # times another equals -3 and adds to -2